OpenStudy (anonymous):

Marilyn knows that she needs $45,000 for a down payment on a house. She found an investment that earns 3.15% interest compounding monthly. She would like to purchase the home in 5 years. How much should she put in the account now to ensure she has her down payment? $38,450.39 $30,871.96 $44,296.89 $52,665.27

OpenStudy (anonymous):

the third choice, i believe.

OpenStudy (phi):

I would use the compound interest formula see http://en.wikipedia.org/wiki/Compound_interest#Compound_Interest \[ \text{Future Value}= P (1+\frac{r}{n})^{nt} \] you are told r is 0.0315, compounded monthly, so n is 12, you are told t is 5 years. you want the future value to be 45000 now solve for P, how much to deposit.

OpenStudy (phi):

if we replace the variables with numbers you get \[ P (1+\frac{0.0315}{12})^{12\cdot 5} = 45000 \] or \[ P (1.002625)^{60}= 45000 \] can you finish ?

OpenStudy (givingtree):

A